### v1

# PAS: Data-Efficient Plug-and-Play Prompt Augmentation System

Abstract.
In recent years, the rise of Large Language Models (LLMs) has spurred a growing demand for plug-and-play AI systems. Among the various AI techniques, prompt engineering stands out as particularly significant. However, users often face challenges in writing prompts due to the steep learning curve and significant time investment, and existing automatic prompt engineering (APE) models can be difficult to use. To address this issue, we propose PAS, an LLM-based plug-and-play APE system. PAS utilizes LLMs trained on high-quality, automatically generated prompt complementary datasets, resulting in exceptional performance. In comprehensive benchmarks, PAS achieves state-of-the-art (SoTA) results compared to previous APE models, with an average improvement of 6.09 points. Moreover, PAS is highly efficient, achieving SoTA performance with only 9000 data points. Additionally, PAS can autonomously generate prompt augmentation data without requiring additional human labor. Its flexibility also allows it to be compatible with all existing LLMs and applicable to a wide range of tasks. PAS excels in human evaluations, underscoring its suitability as a plug-in for users. This combination of high performance, efficiency, and flexibility makes PAS a valuable system for enhancing the usability and effectiveness of LLMs through improved prompt engineering.
Automatic Prompt Augmentation, Plug-and-Play System, Large Language Models, Data Generation, Data Selection.
$\dagger$ The first two authors have equal contributions.
* Corresponding Author
Figure 1. We first present the pipeline of the PAS. PAS takes user prompts, enhances them, and then inputs the augmented prompts into LLMs. As illustrated in the second figure, PAS significantly improves responses across all categories in human evaluation.
1. Introduction
In recent years, rapid advancements in large language models (LLMs)(OpenAI, 2023a; Touvron et al., 2023) have underscored the importance of data management and AI systems in leveraging these technologies(Fernandez et al., 2023; Trummer, 2023; Chen et al., 2023a; Miao et al., 2024; Nie et al., 2023). At the same time, there is an increasing need for scalable plug-and-play systems for LLMs due to their flexibility and efficiency (Al Ridhawi et al., 2020).
Among the various techniques for LLMs, prompt engineering has emerged as a crucial approach due to its extremely low cost and significant enhancement of LLM performance (Sahoo et al., 2024). This technique leverages the inherent capabilities of LLMs to understand and generate human-like text, enabling them to perform a wide range of applications from natural language understanding and generation to specialized domains such as medical diagnosis and legal analysis (Meskó, 2023; Giray, 2023).
In prompt engineering techniques, few-shot learning stands out because it provides a small number of examples to guide the model, thereby enhancing its task-specific performance. Chain-of-Thought (CoT) prompting guides models to reason through problems step-by-step, improving logical consistency and accuracy. In-context learning, by embedding relevant examples and instructions directly within the prompt, allows models to adapt dynamically to new tasks. Despite its potential, prompt engineering is not user-friendly and requires meticulous design. Crafting effective prompts demands a deep understanding of both the model and the task at hand. This process can be time-consuming and often involves extensive trial and error to optimize performance. To tackle those weak points, automatic prompt engineering is designed for easier prompt generation (Zhou et al., 2022a; Pryzant et al., 2023a; Cheng et al., 2023). Although APE models can automatically enhance the prompt, they have to use massive amounts of human-labeled data (Cheng et al., 2023; Ouyang et al., 2022; Rafailov et al., 2024). Additionally, previous methods failed to construct a flexible, user-friendly and effective APE model (Zhou et al., 2022b; Pryzant et al., 2023b). They face the following three key challenges:
C1. Low Efficiency. Previous works on APE primarily use extensive human-labeled instruction datasets, which result in significant human labor (Ouyang et al., 2022; Cheng et al., 2023; Rafailov et al., 2024). Additionally, some methods require training a specific APE model for each LLM, leading to a substantial waste of computational resources (Pryzant et al., 2023a; Yang et al., 2023a).
C2. Low Flexibility. Previous works primarily focus on the performance of APE models, overlooking the importance of flexibility, specifically their model-agnostic and task-agnostic capabilities (Pryzant et al., 2023a; Yang et al., 2023a). Low-flexibility APE systems can lead to computational waste and hinder their application on LLMs (Pryzant et al., 2023a). Additionally, these low-flexibility systems have limited applicability across various scenarios, making them less versatile for diverse use cases and reducing their overall effectiveness in practical applications.
C3. Poor Effectiveness. Although several works aim to automatically enhance the quality of prompts, they often rely on outdated evaluation metrics and do not consistently yield significant improvements across various benchmarks and models (Cheng et al., 2023; Ouyang et al., 2022). Additionally, these models fail to include human evaluations, thus lacking valuable feedback from human users (Cheng et al., 2023; Pryzant et al., 2023a).
Figure 2. Case Study 1, Red text is the complementary prompt generated by PAS. We can see PAS can give complementary prompt to avoid logic traps.
To address these issues, as shown in Figure 1(a), we propose PAS, an automatic prompt-enhancing plug-and-play system. Our approach involves two main phases: high-quality prompt selection and automatic complementary prompt generation, followed by fine-tuning of large language models (LLMs). In the prompt selection phase, we begin by using embedding models to extract features from the prompt data. We then apply clustering algorithms to group and deduplicate similar prompts. Following this, we use LLMs to select high-quality prompts and classify them into various categories. In the automatic complementary prompt generation phase, we employ few-shot learning techniques to generate new prompts. These generated prompts undergo a rigorous selection and regeneration process to ensure their quality. The high-quality generated data is then used to fine-tune the LLMs, forming the core of the PAS system.
The core contributions of this paper are summarized as follows:
  • New Perspective. To the best of our knowledge, this work is the first to construct a curated prompt complementary dataset without human labor. Additionally, we are the first to utilize this dataset to train LLMs to construct the PAS to automatically complement user prompts instead of directly modifying them.
  • New Method. We propose a new method that first utilizes diversity and quality selection for prompt construction. Then, few-shot learning is used for prompt complementary dataset construction. We employ GPT to identify and regenerate incorrectly generated data until it is correct. Instead of using human preference data, we only use the prompt complementary dataset to fine-tune LLMs and achieve SoTA performance.
  • SoTA Performance. (1) High Efficiency. Our PAS model requires only 9000 pairs of prompt complementary data to fine-tune LLMs, making it extremely data-efficient. Additionally, our data generation process is entirely automatic and requires no human labor. Moreover, our PAS model can be integrated into any LLM and solve all tasks, achieving SoTA performance. Therefore, we only need to train one LLM, reducing computational costs. (2) High Flexibility. Our PAS model can be plugged into any LLM, and is model and task agnostic. It achieves SoTA performance across all models and tasks, demonstrating its exceptional flexibility. (3) SoTA Performance on Multiple Benchmarks. Our PAS model achieves SoTA performance across multiple models and comprehensive benchmarks. It also outperforms the previous SoTA model, BPO, under the same experimental settings. Furthermore, it achieves superior performance on human evaluation metrics, highlighting the outstanding capabilities and potential uses of the PAS model.
2. Related Work
2.1. Automatic Prompt Engineering
The effectiveness of large language models in various applications largely depends on the quality of the prompts used. There are already many designed prompts that can significantly enhance the performance of LLMs (Kojima et al., 2022; Wei et al., 2022; Yao et al., 2024; Besta et al., 2024; Yang et al., 2024; Wang et al., 2023). However, these methods that rely on manual prompt engineering are far less scalable. In the field of mathematical logical reasoning for LLMs, the Chain of Thought and its derived strategies are widely popular due to their effectiveness. Zero-shot CoT (Kojima et al., 2022) is adding a simple sentence like “Let’s think step by step” at the end of questions to assist LLMs in generating reasoning steps. Instead of Zero-shot CoT, Manual-Cot (Wei et al., 2022) provides reasoning steps as few shots. Self-Consistency further improves language models’ reasoning performance by generating a diverse set of reasoning paths and choosing the most consistent answer in the final answer set. Tree of Thought (TOT) (Yao et al., 2024) and GOT (Besta et al., 2024) extend the reasoning pathway from linear to non-linear data structures by leveraging multiple LLM queries to elicit different plausible reasoning paths (Yang et al., 2024). Buffer of Thought(BOT) (Yang et al., 2024) designs a series of thought-template for tasks, and for each problem, it retrieve a relevant thought-template to prompt LLMs. PS prompting (Wang et al., 2023) improves COT by encouraging LLMs to devise a plan before attempting to solve a problem.
All the aforementioned prompting engineering strategies have been crafted by human expertise. To avoid manual effort, there is a lot of recent work to explore how to conduct automated prompt engineering (Zhang et al., 2022; Shum et al., 2023; Zhou et al., 2022b; Yang et al., 2023a; Pryzant et al., 2023b; Guo et al., 2023; Fernando et al., 2023). Auto-Cot (Zhang et al., 2022) partitions questions of a given dataset into a few clusters and generates reasoning chains to construct demonstrations for each cluster for Few-shot COT. Automatic-COT (Shum et al., 2023) creates rationale chains to augment exemplars and filters out incorrect ones by checking against the ground truth. Both of them improve the performance of vanilla COT (Kojima et al., 2022; Wei et al., 2022). Unlike previous works, OPRO (Yang et al., 2023a), APO (Pryzant et al., 2023b), and APE (Zhou et al., 2022b) provide an optimizer’s perspective for automatically finding prompts. OPRO (Yang et al., 2023a) leverages LLMs as optimizers, using the accuracy of training datasets—unavailable in real-world scenarios—as the objective value. APO (Pryzant et al., 2023b) provides detailed guidance on prompt refinement at each step, based on the differences between responses and targets. Evoprompt (Guo et al., 2023) and Promptbreeder (Fernando et al., 2023) introduce evolutionary algorithms (EAs) into discrete prompt optimization for specific domains. Similar to evolutionary algorithms, they require evaluating the fitness of each individual prompt in the population, presenting significant challenges in practical applications. Additionally, exploring dozens of generations of prompts imposes a considerable burden.
2.2. Plug-and-Play Systems
Plug-and-play systems have garnered significant attention in recent years due to their modularity and ease of integration in various machine-learning workflows. These systems are designed to operate seamlessly with existing frameworks, allowing for quick and flexible augmentation of functionalities without the need for extensive reconfiguration (Abdulrazak and Helal, 2006; Zhang et al., 2021; Venkatakrishnan et al., 2013).
In image processing research, plug-and-play systems are commonly applied for its outstanding flexibility. Image reconstruction, denoising, deblurring, image enhancement, and super-resolution are all fields where plug-and-play systems are highly needed. By integrating various image processing modules into a unified framework, plug-and-play systems can flexibly combine different methods to achieve better image processing results. Moreover, this system allows for the easy addition or replacement of new processing modules without the need to redesign the entire algorithm, significantly improving the efficiency and effectiveness of image processing.
In the field of artificial intelligence, the rapid advancement of machine learning models has spurred a growing demand for plug-and-play systems. These systems enable seamless integration and adaptation of AI technologies across various applications. Al Ridhawi et al. (2020) have underscored the critical role of AI plug-and-play systems in enhancing scalability, flexibility, and usability in modern computational frameworks.
2.3. Data Quality and Data Selection
The advent of large language models has brought about a substantial increase in the volume of training data (Touvron et al., 2023; OpenAI, 2023b). In this scenario, the quality of data becomes paramount. LLMs, trained on vast amounts of data, can capture subtle nuances and complex patterns in language, excelling in various natural language processing tasks. However, the increase in data volume also brings new challenges, particularly in data quality and data selection (Bai et al., 2024). In this section, we mainly discuss the effectiveness of data quality and data selection.
Data Quality
: High-quality data can significantly enhance the performance of models. (meta llama, 2024) As the volume of data increases, ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation. (Bai et al., 2024) Poor quality data can lead to models learning incorrect patterns and making inaccurate predictions.
Data Selection
: LLMs-based methods were commonly used in data selection. (Bai et al., 2024) For instance, Du et al. (2023) leverages DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023b) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster.
Figure 3. Pipeline for selecting prompt data and generating complementary prompt data.
3. Method
In this section, we first summarize the collection and process of prompt data in section 3.1. Then we introduce the prompt complementary data generation pipeline in section 3.2 to automatically generate high-quality prompt complementary data. After the prompt dataset is generated, we give a comprehensive analysis of the dataset in section 3.3. At last, in section 3.4, we introduce utilizing the prompt augmentation dataset for LLM fine-tuning and then use the fine-tuned LLM to construct a plug-and-play APE system.
3.1. Prompts Data Collection
In this section, we introduce the prompt collection process. To generate high-quality (prompt, complementary prompt) pairs, we first need to select high-quality prompts. To achieve this, we use two curated datasets: the LMSYS-1M dataset (Zheng et al., 2023) and the WildChat dataset (Zhao et al., 2024). As shown in Figure 3(a), our data selection process involves three main steps to ensure the quality and relevance of the data. First, we perform deduplication using the SimCSE bge model to embed the prompts, followed by the HNSW clustering algorithm to group these embeddings. From each cluster, we extract a small amount of data to reduce redundancy. Second, we conduct quality filtering using the BaiChuan 13b model to score the data and filter out low-quality entries, thereby enhancing the overall data quality. Finally, we perform classification. We utilize 60,000 internally labeled classification data points from BaiChuan to fine-tune (SFT) a BaiChuan 13b (Yang et al., 2023b) model, resulting in a classification model that we then use to categorize the prompts into commonly used categories such as Q&A and coding. This series of steps ensures the diversity, quality, and accurate classification of the data.
3.2. Prompts Augmentation Dataset
Input: Golden Data $D_{\text{golden}}=\{(p_{i},a_{i})\}_{i=1}^{N}$, Prompt Set $P=\{p_{j}\}_{j=1}^{M}$
Output: Generated Data $D_{\text{generated}}$
1 $D_{\text{generated}}\leftarrow\emptyset$;
2
3for each $p_{j}\in P$ do
4       $a_{j}\leftarrow\text{FewShotGenerate}(p_{j},D_{\text{golden}})$;
5       $D_{\text{generated}}\leftarrow D_{\text{generated}}\cup\{(p_{j},a_{j})\}$;
6
7for each $(p_{j},a_{j})\in D_{\text{generated}}$ do
8       if not $\text{IsCorrectPair}(p_{j},a_{j},D_{\text{golden}})$ then
9             $D_{\text{generated}}\leftarrow(D_{\text{generated}}-\{(p_{j},a_{j})\})$;
10             while not $\text{IsCorrectPair}(p_{j},a_{j},D_{\text{golden}})$ do
11                   $a_{j}\leftarrow\text{FewShotGenerate}(p_{j},D_{\text{golden}})$;
12
13            $D_{\text{generated}}\leftarrow D_{\text{generated}}\cup\{(p_{j},a_{j})\}$;
14
15
16Function FewShotGenerate($p,D_{\text{golden}}$):
Input: Prompt $p$, Golden Data $D_{\text{golden}}$
Output: Generated Answer $a$
17 $a\leftarrow\text{FewShot Learning with }D_{\text{golden}}$;
18 return $a$;
19
20Function IsCorrectPair($p,a,D_{\text{golden}}$):
Input: Prompt $p$, Answer $a$, Golden Data $D_{\text{golden}}$
Output: Boolean $isCorrect$
21 $isCorrect\leftarrow\text{FewShot Eval with }D_{\text{golden}}$;
22 return $isCorrect$;
23
Algorithm 1 Prompt Augmentation Dataset Generation
To generate a high-quality prompt complementary dataset, we designed an automated data generation pipeline based on Few-Shot Learning. The algorithm mainly consists of two phases: data generation and data selection with regeneration. We use a set of golden data $D_{\text{golden}}=\{(p_{i},a_{i})\}_{i=1}^{N}$, containing 4 to 5 pairs of few-shot examples for each category from BaiChuan. Then, we utilize the prompt dataset $P$ from Section 3.1 to generate high-quality (prompt, complementary prompt) pairs.
Data Generation
For each prompt $p_{j}\in P$ in every category, we utilize the Few-Shot Learning method based on the prompt in Figure 4 to generate a corresponding complementary prompt $a_{j}$ based on the golden data $D_{\text{golden}}$. The generated prompt-complementary prompt pair $(p_{j},a_{j})$ is then added to $D_{\text{generated}}$.
Data Selection and Regeneration
We observed that not all the generated complementary prompt data are of high quality or useful for the original prompt. To address this issue, we proposed a data selection and regeneration pipeline for high-quality complementary prompts.
For each generated prompt-answer pair $(p_{j},a_{j})\in D_{\text{generated}}$, we use Few-Shot Learning based on the prompt in Figure 5 to evaluate its correctness. If the evaluation result is incorrect, we remove the pair and use Few-Shot Learning in the data generation phase to regenerate the answer until the correct answer is generated. Finally, we add the correct prompt-answer pair back to $D_{\text{generated}}$.
Through this data selection and regeneration process, we can automatically generate a prompt complementary dataset while ensuring data quality. This process provides reliable data support for subsequent model training. To better summarize the contents of this section, we translated the data selection process into an algorithm and a pipeline. The data generation pipeline is summarized in Figure 3(b) and Algorithm 1.
Figure 4. Complementary Dataset Generation Prompt
Figure 5. Data Selection and Regeneration Prompt
3.3. Prompt Complementary Dataset
Figure 6. Prompt Complementary Dataset Distribution
In this section, we provide a detailed analysis of the generated prompt-complementary dataset. The prompt-complementary dataset consists of approximately 9,000 high-quality (prompt, complementary prompt) pairs. We summarize the dataset distribution in Figure 7. From Figure 7, it is evident that the data are categorized into 14 categories, with each category containing about 500 data points. This broad coverage across various prompt cases demonstrates the dataset’s exceptional generalization ability. Additionally, there is a substantial amount of Coding and Q&A data. Given that Coding and Q&A are among the most commonly used functions, this distribution is reasonable and emphasizes the most critical functionalities.
Despite these advantages in data distribution, our automatic complementary prompt data generation process described in Section 3.2 allows us to control the categories of the generated data. This flexibility enables our method to cater to both general-purpose models and to generate specialized data to enhance prompt capabilities in specific domains.
3.4. Automatic Prompt Complementary Plug-and-Play System PAS
In this section, we first use the generated prompt-complementary dataset in section 3.2 to fine-tune LLMs for prompt-complementary tasks. Given the flexibility of LLMs as automatic prompt-complementary tools that can be integrated into other generative LLMs, these prompt-complementary models can serve as an automatic, plug-and-play system to enhance LLM performance.
Fine-tune LLMs for Prompt Complementary Models
By utilizing the high-quality generated dataset $D_{\text{generated}}$ in section 3.2, we can automatically fine-tune LLMs to have a prompt complementary ability. We use $M_{p}$ to denote an automatically prompt complementary model, and $M$ to denote an LLM, which can be written as the following mathematic formula:
$M_{p}\leftarrow\text{SFT}(M;D_{\text{generated}})$
We call this $M_{e}$ model PAS, a prompt complementary model with can serve as an automatic, plug-and-play system to enhance LLM performance.
PAS Enhances LLMs Performance
For a prompt $p$ from the user, the complementary prompt $p_{c}$ can be generated using the following formula:
$p_{c}=M_{p}(p)$
The enhanced response $r_{e}$ is then given by:
$r_{e}=\text{LLMs}(\text{cat}(p,p_{c}))$
By generating a complementary prompt, the PAS can improve the user’s prompt without altering the original input. As demonstrated in Section 4, PAS effectively enhances the performance of LLMs.
Plug-and-Play LLMs Promoting System
PAS can be plugged into any other LLMs available via public APIs (OpenAI, 2023a) and can also be integrated into models with open parameters (Bai et al., 2023; Touvron et al., 2023). This flexibility allows for a wide range of applications and improvements across different platforms and systems.
The primary advantage of such a system is its ability to seamlessly enhance the capabilities of existing LLMs without the need for extensive retraining or modification. By simply augmenting the input prompts, PAS leverages the strengths of the underlying models while providing a mechanism to improve their performance. This approach is both cost-effective and efficient, enabling better utilization of computational resources and accelerating the deployment of enhanced language models in various applications.
4. Experiments
In this section, we first introduce the experimental setups. We then aim to answer the following questions to verify the effectiveness, efficiency, and robustness of our proposed KeyframeLLM: Q1: Can our PAS achieve SoTA performance compared to previous SoTA methods? Q2: Can our PAS outperform the previous SoTA model with the same base model? Q3: How efficient and flexible is our model compared to previous APE models? Q4: Can PAS achieve SoTA performance in human evaluation, making it a user friendly system? Q5: Can we visualize the advantages of our method? Q6: Do we need data quality selection and regenerate module in our data generation pipeline?
4.1. Experiments Setting
Datasets.
We followed the steps in section 3 and generated a dataset of 9000 high-quality pairs (prompt, complementary prompt).
Models.
For PAS models, we select several smaller models to efficiently train a PAS model. We select Qwen2-7b-Chat (Bai et al., 2023), LLaMA-2-7b-instruct (Touvron et al., 2023). Then we utilize our trained PAS models to some massive SoTA models, i.e. GPT-4-turbo-2024-04-09, GPT-4-1106-preview, GPT-4-0613, GPT-3.5-turbo-1106 (OpenAI, 2023a), Qwen2-72B-Chat (Bai et al., 2023), and LLaMA-3-70b-instruct (Touvron et al., 2023; meta llama, 2024).
Baselines.
We compare the performance of PAS with models without PAS. Additionally, we compare the performance of PAS with the previous SoTA automatic prompt engineering method BPO (Cheng et al., 2023) to demonstrate the effectiveness of PAS.
Evaluation.
To evaluate the effectiveness of our PAS model, we used three comprehensive benchmarks to thoroughly assess the model’s performance:
  • Arena-hard: This benchmark is designed to test the robustness of language models in handling complex and challenging scenarios. It includes tasks that require advanced reasoning, problem-solving, and understanding of nuanced language constructs. Models are evaluated based on their ability to navigate these hard scenarios and provide accurate, coherent responses.
  • Alpaca-Eval 2.0: This benchmark assesses the general performance of language models across a wide range of standard tasks. It includes a variety of question types and subject areas, ensuring a comprehensive evaluation of the model’s capabilities. The Alpaca-Eval 2.0 is a standard for measuring the overall effectiveness and versatility of language models.
  • Alpaca-Eval 2.0 (LC): This is a variation of the Alpaca-Eval 2.0 benchmark, specifically focusing on language comprehension (LC). It includes tasks that test the model’s ability to understand and interpret complex language inputs accurately. This benchmark is crucial for evaluating the depth of language understanding and the precision of responses generated by the model.
Settings.
For Qwen2-7B-Chat (Bai et al., 2023), LLaMA-2-7B-Instruct (Touvron et al., 2023), Qwen2-72B-Chat (Bai et al., 2023), LLaMA-2-7B-Instruct (Touvron et al., 2023), and LLaMA-3-70B-Instruct (Touvron et al., 2023; meta llama, 2024), we primarily use the hyperparameters from the official repositories. For the GPT model series, we access the models via API. All experiments are conducted on a machine equipped with 8 NVIDIA H100 GPUs, a 120-core CPU, and 960GB of memory.
Table 1. Comparison of PAS, BPO and not using APE (baseline). We can see PAS significantly outperform the baseline with an average improvement of 8 points. Additionally, when compared to the previous state-of-the-art model, BPO, our model still exceeds it by an average of 5.82 points.
Main Model APE-model Arena-hard Alpaca-Eval 2.0 Alpaca-Eval 2.0 (LC) Average
GPT-4-turbo-2024-04-09 None 76.6 46.12 55.02 59.25
GPT-4-1106-preview None 74.8 50 50 58.27
GPT-4-0613 None 37.9 15.8 30.2 27.97
GPT-3.5-turbo-1106 None 18.9 9.2 19.3 15.8
Qwen2-72B-Chat None 48.1 31.7 39.24 39.68
LLaMA-3-70b-instruct None 41.1 33.18 34.42 36.23
Average None 49.57 31.0 38.03 39.53
GPT-4-turbo-2024-04-09 BPO 76.60 54.65 55.28 62.18
GPT-4-1106-preview BPO 74.60 55.19 52.91 60.90
GPT-4-0613 BPO 38.60 19.61 34.08 30.76
GPT-3.5-turbo-1106 BPO 15.90 10.25 20.29 15.48
Qwen2-72b-Chat BPO 44.40 31.25 39.02 38.22
LLaMA-3-70b-instruct BPO 45.20 38.92 39.24 41.12
Average BPO 49.22 34.98 40.14 41.44
GPT-4-turbo-2024-04-09 PAS (PAS-None) 76.9 65.31 56.54 66.62 (+7.37)
GPT-4-1106-preview PAS (PAS-None) 78.8 65.92 53.63 66.12 (+7.85)
GPT-4-0613 PAS (PAS-None) 43.9 34.06 40.33 39.43 (+11.46)
GPT-3.5-turbo-1106 PAS (PAS-None) 22.1 15.82 23.31 20.41 (+4.61)
Qwen2-72B-Chat PAS (PAS-None) 52.2 45.53 44.31 47.35 (+7.67)
LLaMA-3-70b-instruct PAS (PAS-None) 50.3 45.01 40.52 45.28 (+9.05)
Average PAS (PAS-None) 54.03 45.37 43.20 47.53 (+8.00)
GPT-4-turbo-2024-04-09 PAS (PAS-BPO) 76.9 65.31 56.54 66.62 (+4.44)
GPT-4-1106-preview PAS (PAS-BPO) 78.8 65.92 53.63 66.12 (+5.22)
GPT-4-0613 PAS (PAS-BPO) 43.9 34.06 40.33 39.43 (+8.67)
GPT-3.5-turbo-1106 PAS (PAS-BPO) 22.1 15.82 23.31 20.41 (+4.93)
Qwen2-72B-Chat PAS (PAS-BPO) 52.2 45.53 44.31 47.35 (+9.13)
LLaMA-3-70b-instruct PAS (PAS-BPO) 50.3 45.01 40.52 45.28 (+4.16)
Average PAS (PAS-BPO) 54.03 45.37 43.20 47.53 (+6.09)
4.2. Main Experiments
To address Q1, we used Qwen2-7B-Chat as the base model due to its outstanding performance. We subsequently used the prompt complementary data to train a PAS model and compared it to both the baseline model without the APE model and the previous state-of-the-art (SoTA) APE model, BPO (Cheng et al., 2023). We integrated our model into multiple commonly used LLMs, including GPT-4-turbo-2024-04-09, GPT-4-1106-preview, GPT-4-0613, GPT-3.5-turbo-1106 (OpenAI, 2023a), Qwen2-72B-Chat (Bai et al., 2023), and LLaMA-3-70b-instruct (Touvron et al., 2023; meta llama, 2024).
The results in Table 1 clearly illustrate the effectiveness of our PAS method across different models. Compared to the baseline without using APE, PAS shows significant improvements in all metrics, resulting in an average improvement of 8 points, demonstrating the benefits of incorporating prompt complementary data. For instance, in the case of GPT-4-0613, PAS improves the average score by 11.46 points compared to the baseline, highlighting its substantial impact.
Moreover, when compared to the previous state-of-the-art model BPO, our model significantly outperforms it, resulting in an average improvement of 6.09 points. Each model achieves more than a 4-point average improvement across all six base models compared to BPO, with a notable increase of 9.13 points for Qwen2-72B-Chat, indicating a substantial improvement.
BPO is unstable and performs worse than the baseline in some cases, such as GPT-3.5-turbo-1106 and Qwen2-72B-Chat, indicating that the previous SoTA model cannot consistently outperform the baseline. Considering our model exceeds the baseline by 8.00 points and BPO by 6.09 points, it further demonstrates the effectiveness and robustness of our PAS model.
Overall, our PAS method not only outperforms the baseline but also consistently surpasses the previous SoTA model BPO, establishing its robustness and effectiveness as a fine-tuning strategy for enhancing prompt-based learning systems. This consistent performance across various LLMs underscores the robustness of PAS and its potential to set new benchmarks in the field.
4.3. Effectiveness of PAS
To address Q2, we fix the base model and compare our PAS method with the previous BPO (Cheng et al., 2023). We use LLaMA-2-7b-instruct, the same base model as BPO, and utilize the generated complementary prompt data to fine-tune LLaMA-2-7b-instruct. We compare our model performance with BPO.
Table 2. Comparison of PAS and BPO using the same base model, LLaMA-2-7b-instruct, shows that PAS significantly outperforms the BPO model across all LLMs.
Main Model Method Arena-hard Alpaca-Eval 2.0 Alpaca-Eval 2.0 (LC) Average
GPT-4-turbo-2024-04-09 BPO 76.60 54.65 55.28 62.18
GPT-4-1106-preview BPO 74.60 55.19 52.91 60.90
GPT-4-0613 BPO 38.60 19.61 34.08 30.76
GPT-3.5-turbo-1106 BPO 15.90 10.25 20.29 15.48
Qwen2-72b-Chat BPO 44.40 31.25 39.02 38.22
LLaMA-3-70b-instruct BPO 45.20 38.92 39.24 41.12
Average BPO 49.22 34.98 40.14 41.44
GPT-4-turbo-2024-04-09 PAS 73.54 62.58 54.03 63.38 (+1.20)
GPT-4-1106-preview PAS 75.52 64.06 53.07 64.22 (+3.32)
GPT-4-0613 PAS 40.13 33.11 36.70 36.65 (+5.89)
GPT-3.5-turbo-1106 PAS 18.02 16.18 23.67 19.29 (+3.81)
Qwen2-72B-Chat PAS 47.91 40.59 39.99 42.83 (+4.61)
LLaMA-3-70b-instruct PAS 46.30 43.17 38.77 42.74 (+1.62)
Average PAS 50.24 43.28 41.04 44.85 (+3.41)
The results in Table 1 clearly demonstrate the effectiveness of our PAS method across different models. Notably, PAS exhibits a marked improvement in performance metrics compared to BPO, exceeding the baseline by 3.41 points on average. This is particularly evident in models like GPT-4-0613, where the average score improvement is as high as 5.89 points. Even in cases where the improvement is smaller, such as Llama3-70b-instruct, PAS still manages to outperform BPO, indicating its robustness and consistency.
Overall, our PAS method consistently improves model performance across various evaluation models and settings, establishing its effectiveness and robustness as a fine-tuning strategy for enhancing prompt-based learning systems.
4.4. Efficiency and Flexibility of PAS
To address Q3, we compare the data usage and flexibility of PAS with several SoTA models. This comparison highlights the efficiency and versatility of PAS.
4.4.1. Data Usage Comparison
Figure 7. Efficiency of PAS Compared to Other Methods
Figure 7 illustrates the data consumption of PAS compared to other methods, including PPO, DPO, and BPO. The data consumption for each method is as follows:
  • PAS: 9,000k
  • BPO: 14,000k
  • PPO: 77,000k
  • DPO: 170,000k
As shown in the figure, PAS significantly reduces data consumption compared to other methods, demonstrating its efficiency in data usage. It is important to note that ProTeGi and OPRO are not task-agnostic, so they are not included in the data efficiency comparison.
We then calculated the efficiency of PAS using the formula:
$\text{Efficiency}=\frac{\text{Consumption}_{\text{Methods}}}{\text{Consumption}_{\text{PAS}}}$
Based on this formula, PAS achieves an efficiency that is 1.56 times higher than BPO, 8.56 times higher than PPO, and 18.89 times higher than DPO, highlighting its superior data efficiency.
4.4.2. Flexibility Comparison
Table 3 compares the need for human labor and the flexibility of PAS with other models, specifically in terms of being LLM-agnostic and task-agnostic. The table shows that PAS is both LLM-agnostic and task-agnostic, demonstrating its superior flexibility.
Table 3. Comparison of the Need for Human Labor and Flexibility of PAS as a Plug-and-Play System
Method No Human Labor LLM-Agnostic Task-Agnostic
PPO (Ouyang et al., 2022)
DPO (Rafailov et al., 2024)
OPRO (Yang et al., 2023a)
ProTeGi (Pryzant et al., 2023a)
BPO (Cheng et al., 2023)
PAS
4.4.3. Summary
The results indicate that PAS is the only method satisfying all three criteria: no need for human labor, LLM-agnostic, and task-agnostic. This underscores PAS’s superior flexibility and efficiency as a plug-and-play system. Unlike other methods that require significant human intervention and have limitations in applicability across different LLMs and tasks, PAS provides a highly versatile and efficient solution.
In contrast, methods like PPO, DPO, OPRO, and ProTeGi require human labor and lack LLM-agnostic capabilities. Although BPO is both LLM-agnostic and task-agnostic, it still requires human labor, necessitating large human-annotated datasets, whereas PAS can scale up automatically. This highlights PAS’s significant advancements in providing an efficient, flexible, and human-labor-free prompt augmentation system.
4.5. Human Evaluation
To address Q4, we conducted a comprehensive evaluation using human evaluators to assess the online performance of our PAS compared to the baseline model without any prompt augmentation. We compared the grade score benchmark (GSB), availability proportion, full mark proportion, and average score.
From Figure 1, we can see that PAS outperforms the baseline model in various scenarios. Specifically, the results show a higher percentage of wins across different categories such as Analytical Judgment, Subjective Suggestion, Subjective Recommendation, Common Sense, Event Query, Entity Query, Industry Knowledge, and Subject Knowledge. For example, PAS achieves 58.6% wins in Analytical Judgment, 64.3% in Subjective Suggestion, and 61.1% in Common Sense, demonstrating its effectiveness in enhancing performance compared to the baseline without PAS.
As shown in Table 4, our PAS consistently outperforms the baseline, achieving significant improvements in the full mark ratio, average score, and availability ratio. These enhancements across all three evaluation metrics in every benchmark demonstrate the effectiveness of our model. The results indicate not only strong performance on evaluation benchmarks but also positive feedback from human evaluators, showcasing a user-friendly model.
Furthermore, the consistent performance enhancements across all benchmarks highlight the generalization ability and robustness of our model, suggesting its applicability in various domains and its broad impact.
Table 4. Performance Comparison of PAS vs. Non-PAS on Human Evaluation Benchmarks. The PAS consistently outperforms the non-PAS approach across various metrics.
Benchmarks Full Mark Proportion Average Score Availability Proportion Full Mark Proportion (PAS) Average Score (PAS) Availability Proportion (PAS)
Analysis and Judgment 24.14% 3.84 91.38% 43.10% (+18.96) 4.21 (+0.37) 94.83% (+3.45)
Subjective Advice 35.71% 3.71 85.71% 42.86% (+7.15) 3.93 (+0.22) 85.71% (+0.00)
Subjective Recommendation 0.00% 2.4 60.00% 0.00% (+0.00) 2.8 (+0.40) 80.00% (+20.00)
Common Sense 5.56% 3.25 77.78% 27.78% (+22.22) 3.72 (+0.47) 80.56% (+2.78)
Event Query 20.00% 3.3 60.00% 30.00% (+10.00) 3.6 (+0.30) 70.00% (+10.00)
Entity Query 7.32% 3.15 68.29% 9.76% (+2.44) 3.34 (+0.19) 75.61% (+7.32)
Industry Knowledge 20.69% 3.49 78.16% 40.23% (+19.54) 3.78 (+0.29) 79.31% (+1.15)
Academic Knowledge 18.52% 3.35 77.78% 29.63% (+11.11) 3.76 (+0.41) 83.33% (+5.55)
Average 16.49% 3.35 77.78% 27.92% (+11.43) 3.76 (+0.41) 83.33% (+5.55)
4.6. Case Study
To address Q5, this section discusses three different types of case studies that demonstrate the effectiveness of the PAS.
Case Study 1
We examine the effectiveness of PAS in the case study presented in Figure 1. The query, ”If there are 10 birds on a tree and one is shot dead, how many birds are on the ground?”, is a logic trap that can easily mislead both humans and large models if no hints are given. Initially, without the assistance of PAS, GPT responded with an incorrect answer.
Our PAS approach significantly improves this prompt by providing a complementary hint to remind the model that there is a logic trap. With the assistance of PAS, the new response not only avoids the trap but also showcases a multi-step logical reasoning process. This refined answer is clear, accurate, and guides the user through the reasoning process, explaining why only one bird would be on the ground.
From Case Study 1, we can see that PAS is capable of assisting LLMs in avoiding logic traps.
Figure 8. Case Study 2: Red text indicates the response is incorrect. Green text stands for inexplicit sentences. Blue text denotes the no sense sentences. Purple text is the complementary prompt generated bu PAS.
Case Study 2
From case study 2 in Figure 8(a), we can see the user asked about how to quickly boil water in ancient times. However, without the assistance of PAS, the answer from LLM faces the following issues:
  • Instruct Following Issue: The question pertained to boiling water quickly, yet the responses did not adhere to the ”quickly” criterion.
  • Incorrect Response: From the red text, it is evident that using solar energy cannot boil water quickly.
  • Incomplete Explanation: From the green text, the use of fire is mentioned, but the methods to control the fire for quick boiling are not explained.
  • Incorrect Response: From the red text, it is evident that using hot stones cannot boil water quickly either.
  • Logical Flaw: From the blue text, if steam is used, the water is already boiling. Although steam has a high temperature, it cools quickly when it encounters a container, thus it cannot boil water quickly either.
  • Lack of Specific Measures: Fire is the only method suggested for quick heating, but without specific measures, its value is limited.
With the assistance of PAS, we can see in Figure 8(b) that the LLM provides a correct response, effectively avoiding these issues. The improved response follows the ”quickly” criterion, accurately excludes ineffective methods such as solar energy and hot stones, and provides a comprehensive explanation of how to control fire to quickly boil water. This demonstrates the effectiveness of PAS in guiding LLMs to deliver precise and relevant answers.
Figure 9. Case Study 3: Red text is the human comments. Purple text is the complementary prompt generated by PAS.
Case Study 3
From Figure 9, we can see that the user inquired about whether blood pressure increases or decreases when the body loses blood. Typically, a user asking this question is looking for more information about blood pressure changes during blood loss and what actions to take in such a situation. However, as shown in Figure 9(a), the initial response is superficial and, despite being correct, lacks detailed information, which often prompts further questions from the user. In contrast, Figure 9(b) demonstrates that with PAS, a comprehensive and detailed analysis is provided from a physiological and medical perspective on the changes in blood pressure. This approach not only satisfies the immediate query but also equips the user with a thorough understanding, potentially reducing the need for follow-up questions. By delivering such detailed and insightful responses, PAS proves its effectiveness in enhancing user experience and satisfaction in information-seeking scenarios.
From Case Study 3, we can see that PAS can provide more comprehensive answers that consider the user’s potential needs, rather than incomplete ones.
To summarize, the benefits of our PAS method are as follows:
  • Enhanced Context Understanding: PAS breaks down the query into comprehensible parts, ensuring each component is addressed logically and contextually, as demonstrated in Case Study 1 where PAS helped identify and avoid a logic trap.
  • Improved Response Relevance: By complementing and focusing on the query’s intent, PAS minimizes irrelevant or nonsensical responses, thus enhancing the relevance and usefulness of the output, as shown in Case Study 2 where PAS provided a correct and relevant response to quickly boiling water.
  • Comprehensive and Clear Responses: The method promotes detailed explanations, ensuring that the response is not only correct but also easy to understand and logically sound, as illustrated in Case Study 3 where PAS provided a thorough analysis of blood pressure changes.
  • Reduction of Ambiguity: PAS clarifies ambiguities by explicitly stating assumptions and focusing on key elements, thereby providing more accurate and reliable answers, which is evident in all three case studies.
These case studies demonstrate that Automatic Prompt Engineering can significantly elevate the quality of AI interactions, making responses more contextually appropriate, logically consistent, and user-friendly.
4.7. Ablation Study
To address Q6, in this section, following Section 4.2, we first train a PAS model and conduct an ablation study by systematically replacing the data selection and data regeneration modules. Subsequently, we train another PAS model solely based on the data selection module. We compare the performance of these two models and summarize the results in Table 5.
From Table 5, it is evident that excluding the combined data selection and regeneration module leads to a significant decline in our model’s performance across all metrics. On average, our model’s performance decreased by 3.8 points, which is a notable reduction. Specifically, there was a decrease of 6.63 points in the GPT-4-1106-preview benchmark. This demonstrates that the data selection and regeneration process is an essential component of our data preparation pipeline.
Overall, the ablation study highlights the critical role of data selection and data regeneration in enhancing model performance. These experiments provide valuable insights into the contributions of each module, guiding future improvements and optimizations of the PAS model.
Table 5. Performance comparison between PAS models trained on a curated dataset and trained on a dataset without data selection.
Main Model PAS-model Arena-hard Alpaca-Eval 2.0 Alpaca-Eval 2.0 (LC) Average
GPT-4-turbo-2024-04-09 PAS 76.9 65.86 57.09 66.62
GPT-4-1106-preview PAS 78.8 65.92 53.63 66.12
GPT-4-0613 PAS 43.9 34.06 40.33 39.43
GPT-3.5-turbo-1106 PAS 22.1 15.82 23.31 20.41
Qwen2-72B-Chat PAS 52.2 45.53 44.31 47.35
LLaMA-3-70b-instruct PAS 50.3 45.01 40.52 45.28
Average PAS 54.03 45.37 43.20 47.53
GPT-4-turbo-2024-04-09 wo selection 75.0 57.97 49.52 60.83 (-5.79)
GPT-4-1106-preview wo selection 72.2 57.91 48.37 59.49 (-6.63)
GPT-4-0613 wo selection 38.7 31.59 36.19 35.49 (-3.94)
GPT-3.5-turbo-1106 wo selection 20.0 15.88 22.86 19.58 (-0.83)
Qwen2-72B-Chat wo selection 48.9 42.79 41.51 44.40 (-2.95)
LLaMA-3-70b-instruct wo selection 46.0 43.24 38.56 42.60 (-2.68)
Average wo selection 50.13 41.56 39.50 43.73 (-3.80)
5. Conclusion
In recent years, with the development of LLMs, prompt engineering has become increasingly important. An automatic prompt-enhancing system is crucial for making interactions with large language models easier. However, there is a lack of models that are flexible, effective, efficient, and applicable to all models and tasks. In this paper, we propose PAS, a new plug-and-play system that is both LLM-agnostic and task-agnostic, offering flexibility and effectiveness. PAS automatically enhances prompts by complementing them. Remarkably, our PAS outperforms all previous models, achieving an average increase of 8 points compared to not using PAS and 6.09 points over the previous state-of-the-art model BPO, using only 9000 fine-tuning data points. Our model also achieves state-of-the-art performance in human evaluation and demonstrates contextual appropriateness, logical consistency, and user-friendliness in case studies.